# Non-equilibrium dynamic hyperuniform states

### v2

@@ III.1 Stable and metastable non-equilibrium hyperuniform fluids
In the previous section, we mentioned that by modifying the noise in the random organization model to conserve the center-of-mass of the system, one can obtain an active hyperuniform dynamic state with $S(k\rightarrow 0)\sim k^{2}$, and this special noise was recently also shown to enhance the quasi-long range translational order in two-dimensional crystals Galliano et al. (2023); Maire and Plati (2024). However, we note that such center-of-mass conserved noise is unphysical, and it can not be realized in any realistic particle system, because of the inertia of particles. Therefore, in Ref. Lei and Ni (2019), as shown in Fig. 8(A), an athermal dynamical model consisting of $N$ reactive hard spheres in $d$-dimensions with mass $m$ and $\sigma$ at dimensionless particle number density $\tilde{\rho}\equiv\rho\sigma^{d}$ was proposed. In this model, particles undergo reciprocal elastic collisions with an additional kinetic energy $\Delta E$ injected during each collision (reaction). Between collisions, particles with velocity $\mathbf{v}_{i}$ follow the equation of motion $m\frac{d\mathbf{v}_{i}(t)}{dt}=-\gamma\mathbf{v}_{i}(t)$, where $\gamma$ is the linear damping coefficient. There are two characteristic length-scales in the system controlling its phase behaviour: (i) the mean-free path $l_{m}\propto\sigma/\tilde{\rho}$ at low density, which represents the typical distance between two subsequent collisions for a particle; (ii) the dissipation length $l_{d}\equiv\sqrt{m\Delta E}/\gamma$, which represents the typical distance that an isolated active particle can travel before stop. It was found that when $l_{d}<l_{m}$, the system eventually gets trapped in an absorbing state with $T_{k}\equiv m\overline{v^{2}}/k_{B}d=0$. On the other hand, when $l_{d}>l_{m}$, the system self-organizes into a homogeneous active fluid state with a positive kinetic temperature $T_{k}>0$. To understand this, the authors considered the balance between the energy injection $W_{\rm driv}\simeq f_{a}\Delta E$ and dissipation $W_{\rm disp}\simeq\bar{v}^{2}/\gamma$ per particle in the system, with the average collision frequency per particle $f_{a}\simeq\bar{v}/2l_{m}$ and $\bar{v}$ is the average particle speed. For the active fluid state in 2D, a driving-dissipation balance $W_{\rm driv}=W_{\rm disp}$ is achieved, and the kinetic temperature could be written as $k_{B}T_{k}/\Delta E=A(l_{d}/l_{m})^{2}/8$ with the prefactor $A$ obtained from data collapsing of simulations. The phase diagram of the system is shown in Fig. 8(B), and the dash curve is the mean-field theory prediction for the phase boundary.
@@
Subsequently, the kinetic pathway of the discontinuous absorbing transition from the non-equilibrium hyperuniform fluid was investigated in Ref. Lei and Ni (2023), which focused on the barrier controlled athermal driven-dissipative hard-sphere fluid at $T=0$ in 2D. The phase diagram is shown in Fig. 10(A), where the orange curve represents the continuous absorbing transition, the red dot marks the tricritical point, and the black dashed curve represents the stability limit of the active state. As shown in Fig. 10(B), the critical hyperuniformity at the continuous absorbing transition with the structure factor scaling $S(k\rightarrow 0)\sim k^{0.45}$ was observed, which is in agreement with the hyperuniform scaling at the CDP universality class critical point. Additionally, they found a new hyperuniform scaling $S(k\rightarrow 0)\sim k^{1.2}$ in the structure of metastable states near the stability limit of the discontinuous absorbing transition, of which the physical origin remains unknown. Furthermore, while studying the discontinuous absorbing transition in the system, the authors found that the metastable dynamic hyperuniform state could be kinetically stable at the thermodynamic limit, as the waiting time for phase transition from the metastable active state to the absorbing state increases with the system size, as shown in the Fig. 10(C). The kinetic pathway of the discontinuous absorbing transition indicates that the transition is triggered by the large-scale fluctuations (i.e. small $k$ correlation in $S(k)$ in Fig. 10(D)) while the local structure (i.e. small $r$ regime in $g(r)$) remains intact (Fig. 10(E)). This suggests that the hyperuniformity found near the absorbing transition stability limit is metastable yet kinetically stable for infinitely large systems, and this challenges the common understanding of metastability in equilibrium, where discontinuous phase transitions can always be triggered by localized fluctuations, i.e., nucleation.
III.2 Hyperuniformity in chiral active particles
@@
Similarly, another chiral active particle system was investigated in Ref Lei et al. (2019) using Brownian dynamics simulations, in which particles interact through a WCA potential and the motion of particles at temperature $T$ is described by the overdamped Langevin equations
@@ III.3 Hyperuniformity in algae systems intermediated by hydrodynamic flows
In Ref. Huang et al. (2021), Huang et al. investigated the system of marine algae cells (Effrenium voratum), which swim circularly at the air–liquid interface, as shown in Fig. 14(A). The authors experimentally prepared samples of interacting swimming cells at the air–liquid interface, which self-organized into a homogeneous steady state, as shown in Fig. 14(C). They obtained the snapshots and the cell distributions, with which they found a hyperuniform structure factor scaling $S(k\rightarrow 0)\sim k^{0.6}$ for various cell number densities, as shown in Fig. 14(D). Moreover, they also performed computer simulations to model the algae cell system, using a regularized Stokeslet model to obtain the fluid flow generated by cell swimming, as shown in Fig. 14(B). Then, they proposed a particle based model to describe the motion of algae cells
@@
where $\overline{\bm{V}}(\mathbf{R};v_{c})$ is the fluid flow from the Stokeslet model for cell circling with velocity $v_{c}$, and particles interact through hydrodynamic flow with a long cutoff length. As represented by the last term in Eq. 21, particles undergo adjacent jumps with random temporal steps and displacements chosen from a Pareto distribution. The authors measured density fluctuations after randomly initializing particle systems, which evolved into a steady state, as shown in the inset of Fig. 14(C). The structure factor agrees with the experimental data, as shown in Fig. 14(D), indicating that the hyperuniform structure observed in the system is induced by the hydrodynamic interactions caused by the circular swimming of algae.
@@ III.4 Hyperuniformity in vortex systems induced by active rotors on membrane
In Ref Oppenheimer et al. (2022), Oppenheimer et al. investigated a system of $N$ point vortices induced by membrane rotors, which are disks rotating due to torque $\tau$ in the plane of the membrane or the surface of the fluid, as shown in Fig. 15(A). For a single vortex in an ideal inviscid fluid surface induced by a rotor, the stream function can be written as $\Psi(r)=\frac{1}{4\left(H_{0}\left(\frac{r}{\lambda}\right)-Y_{0}\left(\frac{r}{\lambda}\right)\right)}$ for $r\ll\lambda=\frac{\eta_{2D}}{2\eta_{3D}}$, where $\lambda$ is the Shaffman-Delbrück length, $\eta_{2D}$ is the 2D viscosity on the membrane, and $\eta_{3D}$ is the viscosity of the outer bulk fluid. $H_{0}$ and $Y_{0}$ are zeroth-order Struve function and Bessel function of the second kind, respectively. The dynamics of $N$ point vortices could be written as the Hamilton’s equation
@@
As the Fourier transform of a single point vortice stream function is $\tilde{\Psi}(q)=1/q(q+\lambda^{-1})$, at the long-range limit $\tilde{\Psi}(q\rightarrow 0)=1/q^{2}$. Then, the Hamiltonian in Eq. 24 could be written as $\mathcal{H}[\rho(\mathbf{r})]=\frac{N\Gamma^{2}}{4\pi}\int\frac{S({\bf q})}{q^{2}}{\rm d}{\bf q}=\frac{N\Gamma^{2}}{2}\int\frac{S(\mathbf{q})}{q}{\rm d}q$ with ${\rm d}{\bf q}=2\pi q{\rm d}q$. Thus, to ensure the convergence of Hamiltonian at large scales, the structure factor should have the scaling $S(q\rightarrow 0)\sim q^{\alpha}$ with $\alpha>0$, i.e., hyperuniformity. Moreover, we note that the exponent $\alpha\approx 1.3$ found here is different from the value of 2 obtained in the theory and simulations of hyperuniform fluids of active spinners without considering hydrodynamic effects Lei and Ni (2019), and this suggests that hydrodynamic effects can indeed change the exponent of the scaling of structure factor in dynamic hyperuniform states of chiral active particles.
@@ IV.2.2 Hyperuniformity in spinodal solid-state dewetting systems
Physically similar to the spinodal decomposition in binary demixing fluids, Salvalaglio et al. experimentally investigated the spinodal solid-state dewetting system by preparing $\text{Si}_{1-x}\text{Ge}_{x}$ based thin films on ultrathin silicon and insulator substrates in a molecular beam epitaxy reactor under ultrahigh vacuum conditions Salvalaglio et al. (2020). Upon annealing at high temperatures, the $\text{Si}_{1-x}\text{Ge}_{x}$ film undergoes a spinodal dewetting, and AFM images of various film thicknesses and annealing temperatures are shown in Fig. 18(A). By analyzing the spectral density from the AFM images, a hyperuniform scaling is observed, as shown in Fig. 18(B), and the hyperuniform scaling observed is different at different temperature during the dewetting process. The authors noted that larger and disconnected structures, such as the fourth sample shown in Fig. 18(A), exhibit the scaling $\psi^{*}(k)\sim k^{4}$ at small $k$, resembling the cases observed in Model-B and Model-H dynamics. In contrast, more connected structures, like those in the first and third samples shown in Fig. 18(A), exhibit the scaling $\psi^{*}(k)\sim k^{6}$ at small $k$. Additionally, the authors also considered the effect of the initial Ge film thickness, comparing spectral density plots as illustrated in Fig. 18(C), while the peak positions are different, both structures exhibit the same scaling $\psi^{*}(k)\sim k^{6}$ at small $k$.
@@ IV.2.4 Hyperuniformity in active field theories for phase separation
where $\phi$ is a scalar field, e.g., composition in phase separating binary liquid or density field in liquid-gas phase separation, and the functional term $g_{X}\mathcal{A}_{X}[\phi]$ signifies the non-equilibrium contribution from AFTs, while the other parts of the equation remain the same as in the Cahn-Hilliard equation. The authors studied three AFTs: the Effective Chan-Hilliard (ECH) model with $\mathcal{A}_{\chi}=\nabla\cdot(\phi\nabla\phi)$ Speck et al. (2014), the Active Model B (AMB) with $\mathcal{A}_{\chi}=\nabla^{2}\left[(\nabla\phi)^{2}\right]$ Wittkowski et al. (2014), and the Active Model B+ (AMB+) with $\mathcal{A}_{\chi}=\nabla\cdot\left[\left(\nabla^{2}\phi\right)\nabla\phi\right]$ Tjhung et al. (2018). By numerically solving Eq. 28 for various AFTs with critical quenching in 2D, and normalizing the spectral density from the result with a time-dependent length-scale $k_{1}=\int k\tilde{\psi}(k,t){\rm d}k/\int\tilde{\psi}(k,t){\rm d}k$ in the scaling regime, all three active field models exhibit the same universal hyperuniform scaling $\sim k^{4}$ at $k\rightarrow 0$, as shown in Fig. 20. This suggests that their large-scale structural properties are similar. The authors also considered the fluctuation $\sigma^{2}(R)$ and higher-order correlation moments (skewness and kurtosis) at the length-scale $R$ by coarse-graining $\phi(r)$. They found that the variance for the CH model and all active models decay with the same scaling $\sigma^{2}(R)\sim R^{-3}$ at large length-scale $R\rightarrow\infty$, consistent with the hyperuniform spectral density. However, they also found that the higher order moments (skewness and kurtosis) for these models do not exhibit any universal behavior, indicating activity-dependent higher-order correlations.
@@ V Concluding remarks
Disordered hyperuniformity has been introduced for more than twenty years, and it is known that in equilibrium systems to realize disordered hyperuniform states one has to use delicately designed long range interactions Torquato (2018), which is highly challenging experimentally. However, long range interactions are not necessary in disordered hyperuniform states in non-equilibrium systems, and in this article, we have briefly reviewed the progress on non-equilibrium dynamic hyperuniform states, which has become an emerging research direction on hyperuniformity in the past decade. We first introduced the background of random organization model, which was originally proposed to describe the chaotic dynamics and absorbing phase transition in sheared non-Brownian suspensions. Afterwards, we showed the critical absorbing transition point in random organization model and other CDP class models that were found to be hyperuniform. Moreover, we also mentioned that by adding a center-of-mass conservation to the random organization dynamics, the active state would have a stronger hyperuniformity, which could be understood as the unusually suppressed noise term in the dynamical mean-field equation. However, the center-of-mass conserved noise is unphysical in nature, as all particles have inertia. Therefore, we introduced the non-equilibrium hyperuniform fluid found in driven-dissipative particles with finite inertia in dry systems such as reactive hard-sphere fluids and chiral active fluids, which undergo similar absorbing phase transition as in the random organization, with critical point and active state hyperuniformity. Moreover, we reviewed driven particles in wet systems such as the circularly swimming algae system and driven motor systems, in which long range hydrodynamic interactions induced hyperuniformity. Lastly, we also recapped about some recent literature on hyperuniform structures found in phase separating systems through spinodal decomposition, including numerical systems like the Model-B, Model-H, and several active field model systems, as well as experimental systems like phase-separating DNA nanostar droplets and dewetting Ge-based film systems. All those non-equilibrium systems offer new possibilities for realizing hyperuniform structures experimentally, and pave the way for fabricating hyperuniform functional materials.

### v3

@@ III.1 Stable and metastable non-equilibrium hyperuniform fluids
In the previous section, we mentioned that by modifying the noise in the random organization model to conserve the center-of-mass of the system, one can obtain an active hyperuniform dynamic state with $S(k\rightarrow 0)\sim k^{2}$, and this special noise was recently also shown to enhance the quasi-long range translational order in two-dimensional crystals Galliano et al. (2023); Maire and Plati (2024). However, we note that such center-of-mass conserved noise is unphysical, and it can not be realized in any realistic particle system, because of the inertia of particles. Therefore, in Ref. Lei and Ni (2019), as shown in Fig. 8(A), an athermal dynamical model consisting of $N$ reactive hard spheres in $d$-dimensions with mass $m$ and diameter $\sigma$ at dimensionless particle number density $\tilde{\rho}\equiv\rho\sigma^{d}$ was proposed. In this model, particles undergo reciprocal elastic collisions with an additional kinetic energy $\Delta E$ injected during each collision (reaction). Between collisions, particles with velocity $\mathbf{v}_{i}$ follow the equation of motion $m\frac{d\mathbf{v}_{i}(t)}{dt}=-\gamma\mathbf{v}_{i}(t)$, where $\gamma$ is the linear damping coefficient. There are two characteristic length-scales in the system controlling its phase behaviour: (i) the mean-free path $l_{m}\propto\sigma/\tilde{\rho}$ at low density, which represents the typical distance between two subsequent collisions for a particle; (ii) the dissipation length $l_{d}\equiv\sqrt{m\Delta E}/\gamma$, which represents the typical distance that an isolated active particle can travel before stop. It was found that when $l_{d}<l_{m}$, the system eventually gets trapped in an absorbing state with $T_{k}\equiv m\overline{v^{2}}/k_{B}d=0$. On the other hand, when $l_{d}>l_{m}$, the system self-organizes into a homogeneous active fluid state with a positive kinetic temperature $T_{k}>0$. To understand this, the authors considered the balance between the energy injection $W_{\rm driv}\simeq f_{a}\Delta E$ and dissipation $W_{\rm disp}\simeq\bar{v}^{2}/\gamma$ per particle in the system, with the average collision frequency per particle $f_{a}\simeq\bar{v}/2l_{m}$ and $\bar{v}$ is the average particle speed. For the active fluid state in 2D, a driving-dissipation balance $W_{\rm driv}=W_{\rm disp}$ is achieved, and the kinetic temperature could be written as $k_{B}T_{k}/\Delta E=A(l_{d}/l_{m})^{2}/8$ with the prefactor $A$ obtained from data collapsing of simulations. The phase diagram of the system is shown in Fig. 8(B), and the dash curve is the mean-field theory prediction for the phase boundary.
@@
Subsequently, the kinetic pathway of the discontinuous absorbing transition from the metastable non-equilibrium hyperuniform fluid was investigated in Ref. Lei and Ni (2023), which focused on the barrier controlled athermal driven-dissipative hard-sphere fluid at $T=0$ in 2D. The phase diagram is shown in Fig. 10(A), where the orange curve represents the continuous absorbing transition, the red dot marks the tricritical point, and the black dashed curve represents the stability limit of the active state. As shown in Fig. 10(B), the critical hyperuniformity at the continuous absorbing transition with the structure factor scaling $S(k\rightarrow 0)\sim k^{0.45}$ was observed, which is in agreement with the hyperuniform scaling at the CDP universality class critical point. Additionally, they found a new hyperuniform scaling $S(k\rightarrow 0)\sim k^{1.2}$ in the structure of metastable states near the stability limit of the discontinuous absorbing transition, of which the physical origin remains unknown.
@@
Furthermore, while studying the discontinuous absorbing transition in the system, the authors found that the metastable dynamic hyperuniform state could be kinetically stable at the thermodynamic limit, as the waiting time for phase transition from the metastable active state to the absorbing state increases with the system size, as shown in the Fig. 10(C). The kinetic pathway of the discontinuous absorbing transition indicates that the transition is triggered by the large-scale fluctuations (i.e. small $k$ correlation in $S(k)$ in Fig. 10(D)) while the local structure (i.e. small $r$ regime in $g(r)$) remains intact (Fig. 10(E)). This suggests that the hyperuniformity found near the absorbing transition stability limit is metastable yet kinetically stable for infinitely large systems, and this challenges the common understanding of metastability in equilibrium, where discontinuous phase transitions can always be triggered by localized fluctuations, i.e., nucleation.
III.2 Hyperuniformity in chiral active particles
@@
Similarly, another chiral active particle system was investigated in Ref. Lei et al. (2019) using Brownian dynamics simulations, in which particles interact through a WCA potential and the motion of particles at temperature $T$ is described by the overdamped Langevin equations
@@ III.3 Hyperuniformity in algae systems intermediated by hydrodynamic flows
In Ref. Huang et al. (2021), Huang et al. investigated the system of marine algae cells (Effrenium voratum), which swim circularly at the air–liquid interface, as shown in Fig. 14(A). The authors experimentally prepared samples of interacting swimming cells at the air–liquid interface, which self-organized into a homogeneous steady state, as shown in Fig. 14(C). They obtained the snapshots and the cell distributions, with which they found a hyperuniform structure factor scaling $S(k\rightarrow 0)\sim k^{0.6}$ for various cell number densities, as shown in Fig. 14(D). Moreover, they also performed computer simulations to model the algae cell system, using a regularized Stokeslet model to obtain the fluid flow generated by cell swimming, as shown in Fig. 14(B). Then, they proposed a particle based model to describe the motion of $i$-th algae cells the at time $n\tau$ during a time step $\tau$
@@
where $\overline{\bm{V}}(\mathbf{R};v_{c})$ is the fluid flow from the Stokeslet model for cell circling with velocity $v_{c}$ and distance $\mathbf{R}$, and particles interact through hydrodynamic flow with a long cutoff length. In Eq. 21, the second term on the right-hand side indicates that particles undergo adjacent jumps with random temporal steps and random displacements. Specifically, the Kronecker delta function $\delta(\cdot)$ indicates that the $i$-th particle jumps at time $n\tau$ if $n$ satisfies modulo operation $\bmod(n,p)=s^{(i)}$ with $p$ a constant value, and $s^{(i)}$ the an integer constant between zero and $p-1$, which is randomly assigned to all particles. $\mathbf{\eta}\left(D^{i}\right)$ indicates that the jumping displacements are independently chosen from a normal distribution with deviation $\sqrt{2D^{(i)}p\tau}$, and $D^{(i)}$ is the diffusivity for the $i$-th particle randomly drawn from a Pareto distribution Huang et al. (2021). The authors measured particle position distribution after randomly initialization and wait the system evolving into a steady state, as shown in the inset of Fig. 14(C). The structure factor agrees with the experimental data, as shown in Fig. 14(D), indicating that the hyperuniform structure observed in the system is induced by the long-range hydrodynamic interactions caused by the circular swimming of algae.
@@ III.4 Hyperuniformity in vortex systems induced by active rotors on membrane
In Ref. Oppenheimer et al. (2022), Oppenheimer et al. investigated a system of $N$ point vortices induced by membrane rotors, which are disks rotating due to torque $\tau$ in the plane of the membrane or the surface of the fluid, as shown in Fig. 15(A). For a single vortex in an ideal inviscid fluid surface induced by a rotor, the stream function can be written as $\Psi(r)=\frac{1}{4\left(H_{0}\left(\frac{r}{\lambda}\right)-Y_{0}\left(\frac{r}{\lambda}\right)\right)}$ for $r\ll\lambda=\frac{\eta_{2D}}{2\eta_{3D}}$, where $\lambda$ is the Shaffman-Delbrück length, $\eta_{2D}$ is the 2D viscosity on the membrane, and $\eta_{3D}$ is the viscosity of the outer bulk fluid. $H_{0}$ and $Y_{0}$ are zeroth-order Struve function and Bessel function of the second kind, respectively. The dynamics of $N$ point vortices could be written as the Hamilton’s equation
@@
As the Fourier transform of a single point vortice stream function is $\tilde{\Psi}(q)=1/q(q+\lambda^{-1})$, at the long-range limit $\tilde{\Psi}(q\rightarrow 0)=1/q^{2}$. Then, the Hamiltonian in Eq. 24 could be written as $\mathcal{H}[\rho(\mathbf{r})]=\frac{N\Gamma^{2}}{4\pi}\int\frac{S({\bf q})}{q^{2}}{\rm d}{\bf q}=\frac{N\Gamma^{2}}{2}\int\frac{S(\mathbf{q})}{q}{\rm d}q$ with ${\rm d}{\bf q}=2\pi q{\rm d}q$. Thus, due to the conservation of Hamiltonian, the structure factor should have the scaling $S(q\rightarrow 0)\sim q^{\alpha}$ with $\alpha>0$, i.e., hyperuniformity, which ensures the convergence of Hamiltonian at large length scales. Moreover, we note that the exponent $\alpha\approx 1.3$ found here is different from the value of 2 obtained in the theory and simulations of hyperuniform fluids of active spinners without considering hydrodynamic effects Lei and Ni (2019), and this suggests that hydrodynamic effects can indeed change the exponent of the scaling of structure factor in dynamic hyperuniform states of chiral active particles.
@@ IV.2.2 Hyperuniformity in spinodal solid-state dewetting systems
Physically similar to the spinodal decomposition in binary demixing fluids, Salvalaglio et al. experimentally investigated the spinodal solid-state dewetting system by preparing $\text{Si}_{1-x}\text{Ge}_{x}$ based thin films on ultrathin silicon and insulator substrates in a molecular beam epitaxy reactor under ultrahigh vacuum conditions Salvalaglio et al. (2020). Upon annealing at high temperatures, the $\text{Si}_{1-x}\text{Ge}_{x}$ film undergoes a spinodal dewetting, and AFM images of various film thicknesses and annealing temperatures are shown in Fig. 18(A). By analyzing the spectral density from the AFM images, a hyperuniform scaling is observed, as shown in Fig. 18(B), and the hyperuniform scaling observed is different at different temperature during the dewetting process. The authors noted that dewetting at different annealing conditions may induce hyperuniform structures of different structure factor scalings. For example, the fourth sample shown in Fig. 18(A) at a higher temperature $T=570^{\circ}{\rm C}$, exhibits the scaling $\psi^{*}(k)\sim k^{4}$ at small $k$, resembling the cases observed in Model-B and Model-H dynamics. In contrast, structures at lower temperatures, like those in the first ($T=470^{\circ}{\rm C}$) and third ($T=500^{\circ}{\rm C}$) samples shown in Fig. 18(A), exhibit the scaling $\psi^{*}(k)\sim k^{6}$ at small $k$. Additionally, the authors also considered the effect of the initial Ge film thickness, comparing spectral density plots as illustrated in Fig. 18(C), while the peak positions are different, both structures exhibit the same scaling $\psi^{*}(k)\sim k^{6}$ at small $k$.
@@ IV.2.4 Hyperuniformity in active field theories for phase separation
where $\phi$ is a scalar field, e.g., composition in phase separating binary liquid or density field in liquid-gas phase separation, and the functional term $g_{X}\mathcal{A}_{X}[\phi]$ signifies the non-equilibrium contribution from AFTs, while the other parts of the equation remain the same as in the Cahn-Hilliard equation. The authors studied three AFTs: the Effective Chan-Hilliard (ECH) model with $\mathcal{A}_{\chi}=\nabla\cdot(\phi\nabla\phi)$ Speck et al. (2014), the Active Model B (AMB) with $\mathcal{A}_{\chi}=\nabla^{2}\left[(\nabla\phi)^{2}\right]$ Wittkowski et al. (2014), and the Active Model B+ (AMB+) with $\mathcal{A}_{\chi}=\nabla\cdot\left[\left(\nabla^{2}\phi\right)\nabla\phi\right]$ Tjhung et al. (2018). By numerically solving Eq. 28 for various AFTs with critical quenching in 2D, and normalizing the spectral density from the result with a time-dependent length-scale $k_{1}=2\pi/l_{1}$, where $l_{1}$ characterizes the cluster length-scales in the scaling regime, all three active field models exhibit the same universal hyperuniform scaling $\sim k^{4}$ at $k\rightarrow 0$, as shown in Fig. 20. This suggests that their large-scale structural properties are similar. The authors also considered the fluctuation $\sigma^{2}(R)$ and higher-order correlation moments (skewness and kurtosis) at the length-scale $R$ by coarse-graining $\phi(r)$. They found that the variance for the CH model and all active models decay with the same scaling $\sigma^{2}(R)\sim R^{-3}$ at large length-scale $R\rightarrow\infty$, consistent with the hyperuniform spectral density. However, they also found that the higher order moments (skewness and kurtosis) for these models do not exhibit any universal behavior, indicating activity-dependent higher-order correlations.
@@ V Concluding remarks
Disordered hyperuniformity has been introduced for more than twenty years, and it is known that in equilibrium systems to realize disordered hyperuniform states one has to use delicately designed long range interactions Torquato (2018), which is highly challenging experimentally. However, long range interactions are not necessary in disordered hyperuniform states in non-equilibrium systems, and in this article, we have briefly reviewed the progress on non-equilibrium dynamic hyperuniform states, which has become an emerging research direction on hyperuniformity in the past decade. We first introduced the background of random organization model, which was originally proposed to describe the chaotic dynamics and absorbing phase transition in sheared non-Brownian suspensions. Afterwards, we showed the critical absorbing transition point in random organization model and other CDP class models that were found to be hyperuniform. Moreover, we also mentioned that by adding a center-of-mass conservation to the random organization dynamics, the active state would have a stronger hyperuniformity, which could be understood as the unusually suppressed noise term in the dynamical mean-field equation. However, the center-of-mass conserved noise is unphysical in nature, as all particles have inertia. Therefore, we introduced the non-equilibrium hyperuniform fluid found in driven-dissipative particles with finite inertia in dry systems such as reactive hard-sphere fluids and chiral active fluids, which undergo similar absorbing phase transition as in the random organization, with critical point and active state hyperuniformity. Moreover, we reviewed driven particles in wet systems such as the circularly swimming algae system and driven motor systems, in which long range hydrodynamic interactions induced hyperuniformity. Lastly, we also recapped about some recent literature on hyperuniform structures found in phase separating systems through spinodal decomposition, including numerical systems like the Model-B, Model-H, and several active field model systems, as well as experimental systems like phase-separating DNA nanostar droplets and dewetting Ge-based film systems.
All those non-equilibrium systems offer new possibilities for realizing hyperuniform structures experimentally, and pave the way for fabricating hyperuniform functional materials. Future research may consider finding hyperuniform structures in other systems that share similar characteristics with those mentioned in this review. For example, just as hyperuniform critical states were found in various systems with CDP class absorbing phase transitions, it would be interesting to investigate systems with similar non-equilibrium transitions, such as general reversible-irreversible transitions Reichhardt et al. (2023), yielding transitions Jocteur et al. (2024), depinning transitions Reichhardt and Reichhardt (2009); Wiese (2024), and even transitions involving time crystals Yao et al. (2020), etc. Additionally, as discussed in Section II.C and III, noises such as the center-of-mass conserved (CMC) noise or the momentum-conserved reciprocal excitations are known to be the origin of hyperuniform structures with $S(k\rightarrow 0)\sim k^{2}$ in various driven-dissipative steady states Hexner and Levine (2017); Lei et al. (2019); Lei and Ni (2019). This idea of manipulating the noise form has recently been generalized in Model-A and Model-B field theories Ikeda (2023a), which appears to be a promising approach for designing hyperuniform systems. Furthermore, it is noteworthy that CMC noise was found to be related to the breaking of the Mermin-Wagner theorem in low dimensions, allowing the existence of 1D or 2D crystals Galliano et al. (2023); Ikeda (2023b, a); Kuroda et al. (2024). A recent study discussed hyperuniformity and center-of-mass conservation in Cahn-Hilliard dynamics, suggesting that a stronger hyperuniform $k^{4}$ scaling for $k\rightarrow 0$ is enhanced by the unique CMC noise, other than $k^{2}$ scaling if there is conventional thermal noise  De Luca et al. (2024). Although most of the systems reviewed here are classical, dynamic hyperuniform states are also expected in quantum systems. For example, the vortex systems in super-conducting materials Le Thien et al. (2017); Rumi et al. (2019); Llorens et al. (2020); Sánchez et al. (2023), as well as boundary-driven open quantum chains Carollo et al. (2017) were found to have a dynamic hyperuniform state Jack et al. (2015). In addition, designing “temporal hyperuniform” systems rather than the spatial ones by performing time series analysis may also be an interesting future direction, which could be used for exploring hidden temporal orders in networks, economics, or social science data.